# ET-Former: Efficient Triplane Deformable Attention for 3D Semantic Scene Completion From Monocular Camera

### v1

@@ Abstract
We introduce ET-Former, a novel end-to-end algorithm for semantic scene completion using a single monocular camera. Our approach generates a semantic occupancy map from single RGB observation while simultaneously providing uncertainty estimates for semantic predictions. By designing a triplane-based deformable attention mechanism, our approach improves geometric understanding of the scene than other SOTA approaches and reduces noise in semantic predictions. Additionally, through the use of a Conditional Variational AutoEncoder (CVAE), we estimate the uncertainties of these predictions. The generated semantic and uncertainty maps will aid in the formulation of navigation strategies that facilitate safe and permissible decision-making in the future. Evaluated on the Semantic-KITTI dataset, ET-Former achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU, while reducing GPU memory usage of existing methods by $25\%-50.5\%$.
@@ I Introduction
Main Contributions: In this work, we address the challenge of accurate semantic estimation by proposing a triplane-based deformable attention model that uses deformable attention mechanism [zhu2020deformable, xia2022vision] to process voxel and RGB features in three orthogonal views. Our approach provides a better geometric understanding from three views than only mono-cam frame and our invented deformable attention also reduces the computational cost of feature processing. To predict occluded and out-of-view regions, we employ a Conditional Variational AutoEncoder (CVAE) to generate semantic voxel predictions and also estimate the uncertainties of the predictions. Our approach takes advantage of the two stages idea from VoxFormer [li2023voxformer], where the first stage estimates occupancy from the mono-camera, providing guidance to the second stage for completing the 3D semantic occupancy map. Moreover, our triplane-based feature processing mechanism improves upon the deformable attention model in VoxFormer by smoothing estimation noise generated by sparse query voxels, and our CVAE generator performs more accurate prediction than other approaches. Our main contributions are as follows:
@@
    We introduce a triplane deformable model that converts 2D image features into a 3D semantic occupancy map by projecting 3D voxel queries onto three orthogonal planes, where self- / cross-deformable attentions are applied. This model significantly reduces GPU memory usage of existing methods by $25\%-50.5\%$ while refining visual and geometric information from multiple views. It achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU on the Semantic KITTI dataset.
@@ II Related Works
Semantic Scene Completion: Semantic Scene Completion (SSC) involves estimating both the geometry and semantic labels of a 3D scene, a task first introduced by SSCNet [li2023sscnet]. Early methods [chen20203d, zhou2022cross, song2017semantic, reading2021categorical] utilized geometric information to correlate 2D image features with 3D space. MonoScene [cao2022monoscene] and LMSCNet [roldao2020lmscnet] with monocular cameras proposed end-to-end solutions that convert RGB images into 3D semantic occupancy maps through convolutional neural networks (CNNs) or U-Net architectures. TPVFormer [huang2023tri] applied triplane features for pixel processing with multiple cameras but struggled with filling in missing areas effectively and heavy memory usages. More recent approaches like OccFormer [zhang2023occformer] and VoxFormer [li2023voxformer] have utilized transformers to enhance feature processing with monocular cameras. VoxFormer decouples occupancy and semantic predictions into two stages to improve accuracy in both semantic estimation and occlusion prediction. Building on VoxFormer, MonoOcc [zheng2024monoocc] and Symphonize [jiang2024symphonize] introduced more complex structures for both stages, further enhancing occupancy prediction and semantic estimation, though at the expense of higher memory usage. However, these methods do not estimate the certainty of semantic predictions, limiting their effectiveness in downstream tasks. To address this, we design a triplane-based deformable attention model to improve the estimation accuracy and incorporate Conditional Variational AutoEncoder (CVAE) method to predict the uncertainty of semantic estimations.
Figure 2: Architecture: We present a two-stage pipeline for processing mono-cam images and generate both a semantic occupancy map and its corresponding uncertainty map. In the second stage, we introduce an efficient triplane-based deformable attention mechanism, which reduces high-dimensional 3D feature processing to 2D computations. To estimate the uncertainty in the semantic map, we incorporate a CVAE method. In the first stage, we also apply the triplane-based deformable attention model to generate the occupancy queries, as shown in Figure 4.
@@ III-A Problem Definition
We formulate the Semantic Scene Completion (SSC) task as a generative problem. Given a monocular image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, where $(H,W)$ represents the image size, our model aims to generate a semantic occupancy map $\mathbf{M}\in\mathbb{R}^{L\times V\times D\times N}$, where $(L,V,D)$ represents the dimensions of the 3D occupancy map and $N$ is the number of semantic classes. The model follows a CVAE formulation:
@@
where $\mathbf{m}_{s}\in\mathbf{M}$ is the generated semantic occupancy map and $\mathbf{c}$ indicates the condition, which contains image features in our approach. The embedding vector $\mathbf{z}$ is encoded by the encoder $p_{\theta}()$, given the image instance $\mathbf{i}\in\mathbf{I}$. As illustrated in Figure 2, the model consists of a two-stage architecture for the SSC task. In the first stage, the occupancy map (queries) $\mathbf{m}_{o}\in\mathbb{R}^{L\times V\times D\times 2}$ is estimated. It serves as the inference (occupancy) for semantic prediction. In the second stage, the semantic occupancy map is generated using the CVAE formulation, with the inferred voxels from stage one as input and conditioned on the RGB image.
Our encoder in the second stage employs ResNet-50 [he2016deep] and FPN [lin2017feature], which have demonstrated high potential in image feature extraction for vision tasks [li2023voxformer, zheng2024monoocc, zhang2022dino], to process the RGB image into a lower-dimensional vector $\mathbf{f}_{r}\in\mathbb{R}^{w,h,d}$ as image features, where $w<W$, $h<H$, and $d$ denotes the feature dimension. Similar to VoxFormer [li2023voxformer], DETR [zhu2020deformable] is used to enhance the occupancy queries $\mathbf{m}_{o}$ by incorporating the image features $\mathbf{f}_{r}$ and outputs an enhanced query features $\mathbf{f}_{o}$ as follows:
$\mathbf{f}_{o}=DETR(\mathbf{z}_{q},\mathbf{p}_{o},\mathbf{f}_{r})=$
@@
where $M$ and $K$ indicate the head number and offset number. $\mathbf{p}_{o}$ represents the 3D position of each query voxel. $\mathbf{z}_{q}=\mathbf{f}_{e}(\mathbf{p}_{o})$ is the positional embedding of the query positions. $\phi_{s}$ scales the normalized position $\hat{\mathbf{p}}_{o}$ into the image feature size $h\times w$, and $\phi_{o}$ calculates the positional offset with $\mathbf{z}_{q}$. $\mathbf{W}_{m}$ and $\mathbf{W}^{\prime}_{m}$ are linear weights. $A_{mqk}=\phi_{a}(\mathbf{z}_{q})$ is also computed from $\mathbf{z}_{q}$, where $\phi_{a}()$ represents Linear layers.
III-B Conditional Variational Autoencoder Formulation
Our approach utilizes a Conditional Variational Autoencoder (CVAE) to generate semantic occupancy maps, leveraging the unique properties of the CVAE method [bao2017cvae, harvey2021conditional, sohn2015learning]: 1. CVAEs enable conditional generation, allowing for the generation of voxels that adhere to specific conditional information. We employ this capability to generate semantic maps that conform to the RGB information from monocular observations. 2. CVAEs learn a latent space conditioned on specific attributes, which is typically smooth and continuous. This property ensures that small changes in the latent space lead to meaningful variations in the generated voxels, enhancing the expressiveness of our semantic maps. 3. CVAEs inherently model uncertainty in the generated data, accommodating cases with multiple valid outputs for the same condition, unlike deterministic models. This probabilistic nature allows our approach to handle uncertainty in monocular semantic occupancy mapping more effectively.
@@ III-B Conditional Variational Autoencoder Formulation
$\mu=\mathbf{l}_{m}(\mathbf{f}_{o}),\;\;\;\log\text{var}=\mathbf{l}_{v}(\mathbf{f}_{o}),$ (3)
@@
where $\mathbf{l}_{m}$ and $\mathbf{l}_{v}$ are linear layers. The variance of the Gaussian distribution represents the uncertainty in the estimation $\mathbf{m}_{u}$, as depicted in Figure 2.
@@
Figure 3: The Second Stage: We propose triplan-based efficient self-deformable attention (ESDA) and efficient cross-deformable attention (ECDA) methods in this stage to generate 3D features. The input for this stage consists of occupancy queries and image features. The query features are projected onto three orthogonal planes based on positional embedding weights. The three planes are filled by the ESDA method and then enhanced with image features, $\mathbf{f}_{r}$, by the ECDA method. Finally, the plane features are concatenated and upsampled to generate the semantic occupancy map.
Figure 4: The First Stage: We apply the novel triplane-based deformable attention method in this stage. The estimated depth pixels are projected to the 3D space as depth queries to the triplane-based deformable model. The output of the first stage is an occupancy map, the occupied voxels in which are then converted to occupancy queries.
III-C Triplane-based 3D Deformable Attention
@@
The triplane-based representation [huang2023tri, shue20233d] encodes 3D features by leveraging three orthogonal 2D feature planes ($XY$, $XZ$, and $YZ$). This approach significantly reduces both memory and computational requirements compared to directly encoding features in dense 3D grids. In comparison to bird’s-eye-view representations, triplane encoding offers a more sufficient representation of the real-world environment. The orthogonality of the three planes provide a robust embedding of spatial features within the 3D space, making it an effective and efficient representation for complex 3D scenes. The deformable attention mechanism [zhu2020deformable, xia2022vision] is efficient in processing high-resolution features with spatial awareness of image pixels. However, current methods [zhu2020deformable, li2023voxformer] still rely on large sets of query and reference points, resulting in high computational costs for 3D feature processing. To address this, we propose an efficient triplane-based deformable attention method that enhances voxel features with image features, improving the efficiency and performance of cross-source feature extraction.
@@
We propose an efficient deformable triplane decoder in stage two. Instead of querying the entire 3D voxel space, we reduce the dimensionality from 3D to 2D by aggregating voxel features onto three orthogonal planes. We then propose efficient self- and cross-deformable attention mechanisms to process the features of these planes. As shown in Figure 3, the query features are aggregated by the positional embedding weights:
$\mathbf{f}_{p}[x,y]=\sum_{z=0}^{Z}\mathbf{p}_{e}(\mathbf{p}_{o})\circ\mathbf{f}_{o},$ (4)
@@
where $x$ and $y$ are the pixel indices in a plane, which also correlate to the $x,y$ in the 3D space, and $z$ represents the third axis in the 3D space. The same aggregation process also applies to other two planes. The notation $\circ$ in Eqn (4) represents element-wise product.
@@
Since queries projected from 3D space to 2D planes only occupy part of the 2D space, there are empty pixels in the 2D space. Therefore, to have a full representation of the 3D space, we must fill in the missing pixel features. We employ the Deformable Transformer [xia2022vision] to complete the missing pixel features, using plane-wide pixel features as queries and sampling reference points at a stride of four pixels. The Efficient Self-Deformable Attention (ESDA) mechanism completes the three planes, as $\mathbf{f}_{\text{self}}=\text{ESDA}(\mathbf{f}_{p})$.
@@
As the decoder follows the CVAE format, we leverage visual features to enhance the three orthogonal planes, where the condition is the image feature $\mathbf{f}_{r}$. We introduce a new Efficient 3D Cross Deformable Attention (ECDA) model to enhance the plane features $\mathbf{f}_{\text{self}}$ with the image feature $\mathbf{f}_{r}$. Since the plane features are converted from 3D to 2D, we retain 3D geometric representation by normally sampling reference points $\mathbf{p}_{r}\in\mathbb{R}^{l,v,d}$ from the 3D voxel grid, where $l\ll L$, $v\ll V$, and $d\ll D$. These points are projected to the image feature plane $\mathbb{R}^{w\times h}$, so we have $N_{r}=l\times v\times d$ reference points $\mathbf{p}_{i}\in\mathbb{R}^{l\times v\times d\times 2}$ in the image feature plane. The filled-in triplane features, $\mathbf{f}_{\text{self}}$, serve as queries in ECDA. Because the reference points are normally sampled from the 3D space, the $x,y$ coordinates are correlated to the triplane $x,y$ positions. We can use convolutional layers directly to calculate the offsets of the reference points, as $\Delta\mathbf{f}_{r}=\text{Conv}(\mathbf{f}_{\text{self}})$, where $\Delta\mathbf{f}_{r}\in\mathbb{R}^{l,v,c}$ and $c$ is the feature dimension. We separate the feature dimension into $d$ vectors to represent the offset of $d$ points in the $z$ axis. Finally, we have $\Delta\mathbf{p}_{i}=\mathbf{l}_{r}(\text{rearrange}(\Delta\mathbf{f}_{r}))$. Then a multi-head attention model is used to process the sampled image features with the plane features as in Equation 5
$\mathbf{f}_{\text{cross}}=\text{MultiHead}(\mathbf{f}_{\text{self}},\mathbf{f}_{r}[\mathbf{p}_{i}+\Delta\mathbf{p}_{i}]),$ (5)
@@
where $\mathbf{f}_{\text{self}}$ is the query and $\mathbf{f}_{r}[\mathbf{p}_{i}+\Delta\mathbf{p}_{i}]$ serves as key and value of the multi-head attention model. The $\mathbf{f}_{\text{cross}}$ from three planes would be concatenated together and processed by several linear layers and upsampled to the semantic map, $\mathbf{m}_{s}=\mathbf{l}_{s}(\text{UpSample}(\mathbf{l}_{c}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])))$, where $\mathbf{l}_{s}$ and $\mathbf{l}_{c}$ are all linear layers. $\mathbf{f}_{\text{cross}}^{l}$, $\mathbf{f}_{\text{cross}}^{h}$ and $\mathbf{f}_{\text{cross}}^{d}$ are features of the 2D planes orthogonal to the $l,h,d$ axes.
@@
Symphonize [jiang2024symphonize] 41.44 13.44 55.78 26.77 14.57 0.19 18.76 27.23 15.99 1.44 2.28 9.52 24.50 4.32 28.49 3.19 8.09 0.00 6.18 8.99 5.39
@@
TABLE I: Semantic Scene Completion: Our ET-Former outperforms other approaches, showing at least $15.16\%$ and $18.12\%$ improvements in IoU and mIoU. For specific semantic categories, such as road, building, car, vegetation, trunk, terrain, pole, and traffic-sign, our model provides the best accuracy.
@@
Figure 5: Semantic Qualitative Results: The blue rectangles indicate missing areas where VoxFormer and MonoScene do not have but Ground Truth contains. The red rectangles marked the areas with wrong semantic estimations.

### v2

@@ Abstract
We introduce ET-Former, a novel end-to-end algorithm for semantic scene completion using a single monocular camera. Our approach generates a semantic occupancy map from single RGB observation while simultaneously providing uncertainty estimates for semantic predictions. By designing a triplane-based deformable attention mechanism, our approach improves geometric understanding of the scene than other SOTA approaches and reduces noise in semantic predictions. Additionally, through the use of a Conditional Variational AutoEncoder (CVAE), we estimate the uncertainties of these predictions. The generated semantic and uncertainty maps will help formulate navigation strategies that facilitate safe and permissible decision making in the future. Evaluated on the Semantic-KITTI dataset, ET-Former achieves the highest Intersection over Union (IoU) and mean IoU (mIoU) scores while maintaining the lowest GPU memory usage, surpassing state-of-the-art (SOTA) methods. It improves the SOTA scores of IoU from $44.71$ to $51.49$ and mIoU from $15.04$ to $16.30$ on SeamnticKITTI test, with a notably low training memory consumption of $10.9$ GB, achieving at least a $25\%$ reduction compared to previous methods. Project page: https://github.com/jingGM/ET-Former.git.
@@ I Introduction
Main Contributions: In this work, we address the challenge of accurate semantic estimation by proposing a triplane-based deformable attention model that uses deformable attention mechanism [zhu2020deformable, xia2022vision] to process voxel and RGB features in three orthogonal views. Our approach provides a better geometric understanding from three views than only mono-cam frame and our deformable attention also reduces the computational cost of feature processing. To predict occluded and out-of-view regions, we employ a Conditional Variational AutoEncoder (CVAE) to generate semantic voxel predictions and also estimate the uncertainties of the predictions. Our approach takes advantage of the two stages idea from VoxFormer [li2023voxformer], where stage 1 estimates occupancy from the mono-camera, providing guidance to stage 2 for completing the 3D semantic occupancy map. Moreover, our triplane-based feature processing mechanism improves upon the deformable attention model in VoxFormer by smoothing estimation noise generated by sparse query voxels, and our CVAE generator performs more accurate prediction than other approaches. Our main contributions are as follows:
@@
    We introduce a triplane deformable model that converts 2D image features into a 3D semantic occupancy map by projecting 3D voxel queries onto three orthogonal planes, where self- / cross-deformable attentions are applied. This model significantly reduces GPU memory usage of existing methods by at least $25\%$ while refining visual and geometric information from multiple views. Evaluated on the Semantic KITTI dataset, it improves the SOTA scores of IoU from $44.71$ to $51.49$ and mIoU from $15.04$ to $16.30$, with a notably low GPU memory consumption of $10.9$ GB.
@@ II Related Works
Semantic Scene Completion: Semantic Scene Completion (SSC) involves estimating both the geometry and semantic labels of a 3D scene, a task first introduced by SSCNet [li2023sscnet]. Early methods [chen20203d, zhou2022cross, song2017semantic, reading2021categorical] utilized geometric information to correlate 2D image features with 3D space. MonoScene [cao2022monoscene] and LMSCNet [roldao2020lmscnet] with monocular cameras proposed end-to-end solutions that convert RGB images into 3D semantic occupancy maps through convolutional neural networks (CNNs) or U-Net architectures. TPVFormer [huang2023tri] applied triplane features for pixel processing with multiple cameras but struggled with filling in missing areas effectively and heavy memory usages. More recent approaches like OccFormer [zhang2023occformer] and VoxFormer [li2023voxformer] have utilized transformers to process features of monocular cameras. VoxFormer decouples occupancy and semantic predictions into two stages to improve accuracy in both semantic estimation and occlusion prediction. Building on VoxFormer, MonoOcc [zheng2024monoocc] and Symphonies [jiang2024symphonize] introduced more complex structures for both stages, further enhancing occupancy prediction and semantic estimation, though at the expense of higher memory usage. However, these methods do not estimate the certainty of semantic predictions, limiting their effectiveness in downstream tasks. To address this, we design a triplane-based deformable attention model to improve the estimation accuracy and incorporate Conditional Variational AutoEncoder (CVAE) method to predict the uncertainty of semantic estimations.
Figure 2: Overall Architecture: We present a two-stage pipeline for processing mono-cam images and generate both a semantic occupancy map $\mathbf{m}_{s}$ and its corresponding uncertainty map $\mathbf{m}_{u}$. In stage 1, we introduce a novel triplane-based deformable attention model to generate the occupancy queries $\mathbf{m}_{o}$ from the given mono-cam images, which reduces high-dimensional 3D feature processing to 2D computations. The detailed artchitecture of stage 1 is shown in Figure 6. In stage 2, we employ the efficient triplane-based deformable attention mechanism (detailed in Figure 3) to generate the semantic map, with the inferred voxels from stage 1 as input and conditioned on the RGB image. To estimate the uncertainty in the semantic map, we incorporate a CVAE method, and quantify the uncertainty using the variance of the CVAE latent samples.
@@ III-A Problem Definition
We formulate the Semantic Scene Completion (SSC) task as a generative problem. Given a monocular RGB image $\mathbf{i}\in\mathbb{R}^{H\times W\times 3}$, where $(H,W)$ represents the image size, our model aims to generate a semantic occupancy map $\mathbf{M}\in\mathbb{R}^{L\times V\times D\times N}$, where $(L,V,D)$ represents the dimensions of the 3D occupancy map and $N$ is the number of semantic classes. The model follows a CVAE formulation:
@@
where $\mathbf{m}_{s}\in\mathbf{M}$ is the generated semantic occupancy map given the image instance $\mathbf{i}$. $p(\mathbf{m}_{s}|\mathbf{z},\mathbf{i})$ denotes the conditional distribution of generating the semantic map $\mathbf{m}_{s}$ given the latent sample $\mathbf{z}$ and image $\mathbf{i}$. $p(\mathbf{z})$ is the prior distribution of $\mathbf{z}$. $p_{\theta}(\mathbf{z}|\mathbf{i})$ represents the approximate posterior distribution of the latent representation $\mathbf{z}$, given the image $\mathbf{i}$.
As illustrated in Figure 2, the model employs of a two-stage architecture for the SSC task. In stage 1, the occupancy map (queries) $\mathbf{m}_{o}\in\mathbb{R}^{L\times V\times D\times 2}$ is estimated using the given image. The inferred $\mathbf{m}_{o}$ serves as the input for semantic prediction of stage 2. In stage 2, the semantic occupancy map $\mathbf{m}_{s}$ is generated using the CVAE formulation, with the inferred voxels from stage 1 as input and conditioned on the RGB image. As a byproduct of the CVAE formulation, reparameterizing the latent variables as Gaussian distributions allows for quantification of uncertainty in the occupancy map $\mathbf{m}_{o}$ from stage 1, giving rise to an uncertainty map $\mathbf{m}_{u}$.
One of our major innovations is the Deformable Triplane Decoder, which is primarily designed for stage 2 but also serves as the backbone of stage 1 to provide more visual information for occupancy prediction. Given this interconnected design, we will first describe the methodology of stage 2 in detail. Subsequently, we will elaborate on the sepecifics of stage 1 in Section III-E, highlighting how the Deformable Triplane Decoder is adapted and utilized in this initial stage.
III-B Voxel Encoder
Our voxel encoder in stage 2 employs ResNet-50 [he2016deep] and Feed Forward Network (FFN) [lin2017feature], which have demonstrated high potential in image feature extraction for vision tasks [li2023voxformer, zheng2024monoocc, zhang2022dino], to process the RGB image into lower-dimensional features $\mathbf{f}_{r}\in\mathbb{R}^{w,h,d}$, where $w<W$, $h<H$, and $d$ denotes the feature dimension. DETR [zhu2020deformable] is used to enhance occupancy queries $\mathbf{m}_{o}$ by incorporating the image features $\mathbf{f}_{r}$ and outputs occupancy features $\mathbf{f}_{o}$, which is shown in Equation 2. This process is referred to as voxel deformable cross attention in Figure 2.
$\mathbf{f}_{o}=DETR(\mathbf{z}_{q},\hat{\mathbf{p}}_{o},\mathbf{f}_{r})=$
@@ III-B Voxel Encoder
where $\mathbf{p}_{o}$ represents the 3D position of each query voxel in $\mathbf{m}_{o}$ and $\hat{\mathbf{p}}_{o}$ indicats its normalized position. $\mathbf{z}_{q}=\mathbf{p}_{e}(\mathbf{p}_{o})$ is the positional embedding of the query positions, and $\mathbf{p}_{e}()$ is a positional embedding function. $M$ and $K$ indicate the head number and the offset number. $\phi_{s}$ scales the normalized position $\hat{\mathbf{p}}_{o}$ into the image feature size $h\times w$, and $\phi_{o}$ calculates the positional offset with $\mathbf{z}_{q}$. $\mathbf{f}_{r}(\phi_{s}(\hat{\mathbf{p}}_{o})+\phi_{o}(\mathbf{z}_{q}))$ is the feature at location $(\phi_{s}(\hat{\mathbf{p}}_{o})+\phi_{o}(\mathbf{z}_{q}))$. $\mathbf{W}_{m}$ and $\mathbf{W}^{\prime}_{m}$ are linear weights. $A_{mqk}=\phi_{a}(\mathbf{z}_{q})$ is also computed from $\mathbf{z}_{q}$, where $\phi_{a}()$ represents Linear layers.
III-C Conditional Variational Autoencoder Formulation
Our approach utilizes a Conditional Variational Autoencoder (CVAE) to generate semantic occupancy maps, leveraging the unique properties of the CVAE method [bao2017cvae, harvey2021conditional, sohn2015learning]: 1. CVAEs enable conditional generation, allowing for the generation of voxels that adhere to specific conditional information. We employ this capability to generate semantic maps that conform to the RGB information from monocular observations. 2. CVAEs learn a smooth and continuous latent space. This property ensures that small changes in the latent space lead to meaningful variations in the generated voxels, enhancing the expressiveness of our semantic maps. 3. CVAEs inherently model uncertainty in the generated data, accommodating cases with multiple valid outputs for the same condition, unlike deterministic models. This probabilistic nature allows our approach to handle uncertainty in monocular semantic occupancy mapping more effectively.
@@ III-C Conditional Variational Autoencoder Formulation
$\hat{\mathbf{f}}_{o}\sim\mathcal{N}\{\mu=\mathbf{l}_{m}(\mathbf{f}_{o}),\;\;\;\log\text{var}=\mathbf{l}_{v}(\mathbf{f}_{o})\},$ (3)
@@
where $\mathbf{l}_{m}$ and $\mathbf{l}_{v}$ are linear layers. $\mu$ and $var$ are the mean and variance of the distribution. The variance of the Gaussian distribution represents the uncertainty in the uncertainty map $\mathbf{m}_{u}$, as shown in Figure 2.
@@
Figure 3: (a) Deformable Triplane Decoder of Stage 2: We propose triplane-based efficient self-deformable attention (ESDA) and efficient cross-deformable attention (ECDA) methods in the decoder to generate 3D features. The input to the decoder consists of Gaussian samples $\hat{\mathbf{f}}_{o}$ and image features $\mathbf{f}_{r}$. The Gaussian samples are projected onto three orthogonal planes based on positional embeddings. The three planes features are completed by ESDA and then enhanced with image features, $\mathbf{f}_{r}$, by ECDA. Finally, the plane features are concatenated and upsampled to generate the semantic occupancy map. (b) Detailed Structure of ECDA: ECDA utilizes image features at deformed points $(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})$ and completed plane feature $\mathbf{f}_{\text{self}}$ to conduct cross attention.
III-D Triplane-based 3D Deformable Attention
@@
The triplane-based representation [huang2023tri, shue20233d] encodes 3D features by leveraging three orthogonal 2D feature planes ($XY$, $XZ$, and $YZ$). This approach significantly reduces both memory and computational requirements compared to directly encoding features in dense 3D grids. In comparison to bird’s-eye-view representations, triplane encoding offers a more sufficient representation of the real-world environment. The orthogonality of the three planes provide a robust embedding of spatial features within the 3D space, making it an effective and efficient representation for complex 3D scenes. The deformable attention mechanism [zhu2020deformable, xia2022vision] is efficient in processing high-resolution features with spatial awareness of image pixels. However, current methods [zhu2020deformable, li2023voxformer] still rely on large sets of query and reference points, resulting in high computational costs for 3D feature processing. To address this, we propose an efficient triplane-based deformable attention method that enhances voxel features with image features $\mathbf{f}_{r}$, improving the efficiency and performance of cross-source feature extraction.
@@
We propose an efficient deformable triplane decoder in stage 2. Instead of querying the entire 3D voxel space, we reduce the dimensionality from 3D to 2D by aggregating voxel features onto three orthogonal planes. We then propose efficient self- and cross-deformable attention mechanisms to process the features of these planes. As shown in Figure 3, the query features $\hat{\mathbf{f}}_{o}$ are aggregated by the positional embedding weights:
$\mathbf{f}_{p}(x,y)=\sum_{z=0}^{Z}\mathbf{p}_{e}(\mathbf{p}_{o}(x,y,z))\circ\hat{\mathbf{f}}_{o}(x,y,z),$ (4)
@@
where $x$ and $y$ are the cell indices in a plane, which also correlate to the $x,y$ in the 3D space, and $z$ represents the third axis in the 3D space. $\mathbf{p}_{e}$ is a positional embedding function. $\hat{\mathbf{f}}_{o}$ is a sample from the normal distribution in Equation 3. The notation $\circ$ in Equation 4 represents element-wise product. In short, $\mathbf{f}_{p}$ at location $(x,y)$ is the weighted sum of $\hat{\mathbf{f}}_{o}(x,y,z)$ across all values of $z$. The same aggregation process also applies to other two planes.
@@
Since queries projected from 3D space to 2D planes only occupy part of the 2D space, there are empty cells in the 2D space. Therefore, to have a full representation of the 3D space, we must fill in the missing cell features. As for the SemanicKITTI dataset [behley2019semantickitti], the missing cells are more than half of the plane, so the plane features are sparse. We employ the Deformable Transformer [xia2022vision] to complete the missing cell features, using plane-wide cell features as queries. The reference points are downsampled from the image feature space at a stride of four pixels to strike a balance between computation efficiency and performance. The Efficient Self-Deformable Attention (ESDA) mechanism completes the three planes, as $\mathbf{f}_{\text{self}}=\text{ESDA}(\mathbf{f}_{p})$.
@@
Since the decoder follows the CVAE format, we leverage image features, $\mathbf{f}_{r}$, as a condition to provide different view information and color information, thereby enhancing the features of the three orthogonal planes. We introduce a new Efficient 3D Cross Deformable Attention (ECDA) model to process the plane features $\mathbf{f}_{\text{self}}$ with the image feature $\mathbf{f}_{r}$, which is illustrated in Figure 3. Since the plane features are converted from 3D to 2D, we retain 3D geometric representation by uniformly sampling reference points $\mathbf{p}_{r}\in\mathbb{R}^{l,v,d}$ from the 3D voxel grid, where $l\ll L$, $v\ll V$, and $d\ll D$. These points in the camera’s view, can be projected to the image feature plane $\mathbb{R}^{w\times h}$, so we have $N_{r}=l\times v\times d$ reference points $\mathbf{p}_{i}\in\mathbb{R}^{l\times v\times d\times 2}$ in the image feature plane. The filled-in triplane features, $\mathbf{f}_{\text{self}}$, serve as queries in ECDA. Because the reference points are uniformly sampled from the 3D space, the $x,y$ coordinates are correlated to the triplane $x,y$ positions. We can use convolutional layers directly to calculate the offsets of the reference points, as $\Delta\mathbf{p}_{r}=\text{Conv}(\mathbf{f}_{\text{self}})$, where $\Delta\mathbf{p}_{r}\in\mathbb{R}^{l\times v\times c}$ and $c$ is the feature dimension. We separate the feature dimension into $d$ vectors to represent the offset of $d$ points in the $z$ axis. Finally, we have $\Delta\mathbf{p}_{i}=\mathbf{l}_{r}(s(\Delta\mathbf{p}_{r}))$, where $s()$ represents the feature separation, and $\mathbf{l}_{r}$ is a linear layer. Then a multi-head cross attention model is used to process the sampled image features with the plane features as in Equation 5:
$\mathbf{f}_{\text{cross}}=\text{CrossAttention}(\mathbf{f}_{\text{self}},\mathbf{f}_{r}(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})),$ (5)
@@
where $\mathbf{f}_{\text{self}}$ is the query. $\mathbf{f}_{r}(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})$, the image feature at deformed point locations $(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})$, serves as key and value of the multi-head attention model. The $\mathbf{f}_{\text{cross}}$ from three planes would be concatenated together and processed by several linear layers and upsampled to the semantic map, $\mathbf{m}_{s}=\mathbf{l}_{s}(\text{UpSample}(\mathbf{l}_{c}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])))$, where $\mathbf{l}_{s}$ and $\mathbf{l}_{c}$ are all linear layers. $\mathbf{f}_{\text{cross}}^{l}$, $\mathbf{f}_{\text{cross}}^{h}$ and $\mathbf{f}_{\text{cross}}^{d}$ are features of the 2D planes orthogonal to the $l,h,d$ axes.
@@
Symphonies [jiang2024symphonize] 42.19 15.04 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00
@@
TABLE I: Quantitative Comparison On SemanticKITTI test [behley2019semantickitti]: Our ET-Former outperforms other camera-based approaches, improving the SOTA scores of IoU from $44.71$ to $51.49$ and mIoU from $15.04$ to $16.30$. For specific semantic categories, such as building, car, other-vehicle, vegetation, trunk, terrain, and pole, our model provides the best accuracy.
@@
Figure 4: Semantic Qualitative Results: The blue rectangles highlight areas that are present in the Ground Truth but are missing from both VoxFormer and MonoScene predictions. The red rectangles marked the areas with wrong semantic estimations from Voxformer and MonoScene.